on a scatter plot of data,a trend line passes through the points (-6,7) and (-2,10). describe the correlation.
Answer by stanbon(75887)

on a scatter plot of data,a trend line passes through the points (-6,7) and (-2,10). describe the correlation.
--------------
slope = (10-7)/(2--6) = 3/8
---------------------------------
Form: y = mx + b
You have x,y,and m ; solve for "b":
10 = (3/8)*2 + b
10 = (3/4) + b
b = 37/4
-------------------
Equation: y = (3/8)x + (37/40)
======================================
Cheers,
Stan H.